OpenStudy (anonymous):

2. Find the measure of b, given m∠A = 38°, m∠B = 74°, and a = 31. (1 point)

OpenStudy (tyteen4a03):

Use the sine formula: \(\frac{sin(A)}{31} = \frac{sin(B)}{b}\). (Forgot if it's in the correct place - if not just swap the sines)

OpenStudy (anonymous):

i don't really know how to use them

OpenStudy (tyteen4a03):

@artem151 Might need to go back and revise more then, but here's a dumbed-down version of it:  In here, \(\frac{sin(A)}{a} = \frac{sin(b)}{b} = \frac{sin(C)}{c}\).

OpenStudy (anonymous):

so the answer would be 48.4